OpenStudy (anonymous):

(-3X) + 3 = 18 -5 -3 15 -2

OpenStudy (anonymous):

-3x + 3 = 18 (subtract 3 from both sides) -3x = 18 - 3 -3x = 15 (now divide both sides by -3 to get x by itself) -3x/-3 = 15/-3 x = -5

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no ha thank you! c: